An electron is released from rest at the negative plate of a parallel plate capacitor and accelerates to the positive plate. The plates are separated by a distance of 1.2 cm and the electric field within the capacitor has a magnitude of 2.1 * 10^6 V/m. What is the speed of the electron just as it reaches the positive plate?
Answer #1

Voltage across capacitor

V=Ed =2.1*106*0.012=25200 Volts

By Conservation of energy

(1/2)mv2=qV

(1/2)(9.11*10-31)*v2 =(1.6*10-19)*25200

v=9.41*107 m/s